OpenStudy (anonymous):

Part 1: Show the work to solve |x+3| +7 less than or equal to 10 Part 2: Describe the graph of the solution in a complete sentence.

OpenStudy (anonymous):

\[-6\le x \le 0\]\[\left| x+3 \right|\le3\] This translates to: \[x+3\le3, x+3\ge-3\] Solving for these